OpenStudy (anonymous):

Rewrite the rational exponent as a radical. 5 to the 3 over 4 power, to the 2 over 3 power

OpenStudy (anonymous):

\[\large \sf (5^{\frac{3}{4}})^{\frac{2}{3}}\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

How about \[\large 5^{(\frac{3}{4}^{\frac{2}{3}})}\]

OpenStudy (jdoe0001):

what the?

OpenStudy (jdoe0001):

sounds to me a lot like \(\bf \large (5^{\frac{3}{4}})^{\frac{2}{3}}\)

OpenStudy (anonymous):

That was my first guess /: